OpenStudy (anonymous):

4 ( 2x - 3 ) = 2 ( 3x - 1 )

OpenStudy (jlastino):

Distribute 4 and 2 to (2x-3) and (3x-1) respectively 8x-12 = 6x-2 Then transpose 6x to the left side and -12 to the right 8x-6x=-2+12 2x=10 x=5

OpenStudy (anonymous):

How do you solve this system of equations ? 3x + y = 11 y = x + 3

OpenStudy (jlastino):

Since y=x+3, you can substitute that value of y into the first equation (3x+y=11) This will give 3x+(x+3)=11 4x+3=11 4x=11-3 4x=8 x=2

OpenStudy (anonymous):

in 1978, Franklin Jacobs high-jumped 7 feet 7 1/4 inches. Jacobs was 5 feet 8 inches tall. how many more inches did jacob jump than his height ?